OpenStudy (firejay5):

How do you graph x = 2y^2 - 3?, and give the Domain (x) and Range (y)

OpenStudy (anonymous):

solve for y

OpenStudy (firejay5):

yea

OpenStudy (anonymous):

then move your vertex

OpenStudy (firejay5):

I mean I haven't graphed it yet and needed to be y =

OpenStudy (firejay5):

How would I graph it the way it is or is it possible

OpenStudy (anonymous):

ok tell me what you go when you solved for y. y= and some equation right?

OpenStudy (firejay5):

like explain it to you

OpenStudy (anonymous):

yeah tell me your new equation

OpenStudy (firejay5):

that's what I need help on actually

OpenStudy (anonymous):

just reverse you pemdas.

OpenStudy (firejay5):

how would I do that

OpenStudy (anonymous):

add 3 to both sides first, divide both sides by 2 and square root both sides

OpenStudy (firejay5):

I think it's wrong, but I got: y = 3x + 2^2

OpenStudy (anonymous):

no. \[\sqrt{\frac{ x+3 }{ 2}}=y\]

OpenStudy (firejay5):

y = mx + b

OpenStudy (firejay5):

write it out like that

OpenStudy (anonymous):

yu cant because yur line above has the x and y on the opposite sides. was it suppose to be that way?

OpenStudy (firejay5):

It has to be written out y intercept form

OpenStudy (anonymous):

then i cant help yu ther because i dont think it can.

OpenStudy (anonymous):

alryte sir

OpenStudy (firejay5):

awesome

OpenStudy (anonymous):

to graph this the first thing would be to set it to y= format

OpenStudy (firejay5):

yep I got that it's just how to I forgot to do that

OpenStudy (firejay5):

whenever I get into y = form I am good

OpenStudy (anonymous):

herewego pretty much got it with \[y=\pm \sqrt{\frac{ x+3 }{ 2 }}\]

OpenStudy (firejay5):

did you subtract 3 to both sides

OpenStudy (anonymous):

no i just rearranged it so that 2y^2 was by itself

OpenStudy (firejay5):

was that the answer

OpenStudy (firejay5):

so what's the answer

OpenStudy (anonymous):

to find the domain would be \[\frac{ x+3 }{ 2 }\ge 0\]

OpenStudy (firejay5):

well domain (x) is -2, -1, 0, 1, 2

OpenStudy (firejay5):

I just need range (y)

OpenStudy (anonymous):

range is your y values so it appears to be 0 to infinity

OpenStudy (firejay5):

so would it be 0 for all y values

OpenStudy (firejay5):

uh correct

OpenStudy (anonymous):

\[[0,\infty)\]

OpenStudy (firejay5):

is that it

OpenStudy (firejay5):

uh???

OpenStudy (anonymous):

ya

OpenStudy (firejay5):

why nick???

OpenStudy (anonymous):

well the y value starts at 0 and goes off to infinity

OpenStudy (firejay5):

is it a function

OpenStudy (anonymous):

im basing it off the graph i sent the link

OpenStudy (firejay5):

At least it's right